# The Jaynes Cummings model as an autonomous Maxwell demon

### v1

@@ Abstract
We revisit the Jaynes cummings model as an autonomous thermodynamic machine, where a qubit is coupled to a cavity with a very large initial coherent field. Our analysis reveals different behaviors of this model including the expected unitary behavior of qubit at short times, which fails at longer times when the cavity starts measuring the qubit. The subsequent dynamics feature feedback from the cavity on the qubit, which appears to purify the qubit irrespective of its initial state. Upon analyzing these dynamics from a thermodynamic standpoint, we found that the cavity behaves as an autonomous Maxwell demon.
Autonomous machines, performing tasks without external control, are an important concept in thermodynamics [1, 2]. Their time-dependent evolution derive solely from their initial non-equilibrium condition, akin to automaton. This concept is particularly important in the case of quantum machines [3, 4, 5, 6, 7, 8, 9], where the coherence of microscopic constituents is crucial, as avoiding classical controllers empirically mitigates decoherence. This motivates research into autonomous (or coherent) feedback [10, 11, 12, 13, 14, 15], where a quantum system is able to drive a target system conditionally on the values of a few of its observables, e.g. applied to quantum error correction [16, 17, 18]. When these controllers perform thermodynamic tasks, such as work extraction or cooling, they are called autonomous Maxwell demons [19, 20, 21, 22, 23, 24, 25]. A thorough characterization of these machines is of major fundamental interest, as it is linked to the long-standing question of how the dynamics induced by quantum measurements [26, 27, 28] and the laws of thermodynamics [29, 30, 31] emerge in an entirely quantum framework.
@@
Model – We consider qubit $Q$ (e.g. a two-level atom) interacting with a harmonic oscillator (hereafter dubbed the cavity $C$) according to the Jaynes-Cummings interaction $\hat{V}_{\text{QC}}=i\hbar g(a\sigma_{+}-a^{\dagger}\sigma_{-})$, where $a$ denotes the bosonic annihilation operator of the cavity, and $\sigma_{-}=|g\rangle\langle e|=\sigma_{+}^{\dagger}$ the lowering operator of the atom. We focus on resonant conditions, where the qubit and cavity Hamiltonians are $\hat{H}_{Q}=\hbar\omega_{0}\sigma_{+}\sigma_{-}$ and $\hat{H}_{C}=\hbar\omega_{0}a^{\dagger}a$, respectively. We assume that the cavity is initially prepared in a coherently-displaced thermal state $\hat{\rho}_{C}(0)=D(\alpha_{0})\hat{w}_{\beta(0)}D^{\dagger}(\alpha_{0})$, where $\hat{w}_{\beta(0)}=\frac{e^{-\beta(0)\hbar\omega_{0}a^{\dagger}a}}{Z_{C}(0)}$ is a thermal equilibrium state at inverse temperature $\beta(0)$, with $Z_{C}(0)=\text{Tr}\{e^{-\beta(0)\hbar\omega_{0}a^{\dagger}a}\}$. $D(\alpha_{0})=e^{\alpha_{0}a^{\dagger}-\alpha_{0}^{*}a}$ denotes the phase-space displacement operator with amplitude $\alpha_{0}=\sqrt{n_{0}}e^{i\phi_{0}}$. We are interested in the limit $n_{0}\gg\bar{n},1$, with $\bar{n}=(e^{\beta(0)\hbar\omega_{0}}-1)^{-1}$ the initial thermal occupation of the cavity, in which the cavity is expected to induce a classical drive for the qubit. Due to the resonance condition, the evolution of the total system in the interaction picture is governed by Hamiltonian $\hat{H}_{I}(t)=V_{\text{QC}}$ [32].
Figure 1: (a) Entropy variation $\Delta S_{Q}$ of the qubit over a long time scale for different initial qubit states $|\psi_{1}\rangle=|e\rangle\langle e|$, $|\psi_{2}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+e^{i\frac{\pi}{4}}|g\rangle)$, and $|\psi_{3}\rangle=|+_{y}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+i|g\rangle)$. (b) zoom in the early times of (a), in semilog scale.(c) Excited state population $P_{e}=\langle e|\hat{\rho}_{Q}(t)|e\rangle$ (in blue) and coherence $C_{eg}=\langle e|\hat{\rho}_{Q}(t)|g\rangle$ (in orange and red) of the atom when it is initialized in $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$. The dots correspond to expansions up to second order in $1/n_{0}$ and the solid lines are obtained from exact expressions, truncating the cavity Hilbert space to $N_{\text{ph}}=500$. Parameters: $n_{0}=100$, $\bar{n}=1$.
@@
The autonomous dynamics of the machine features three successive regimes with very different thermodynamic interpretations, that can be identified from the time-evolution of the qubit’s von Neumann entropy as shown in Fig. 1a) and b). We analyze them in the following.
Unitary regime – For times $t\ll g^{-1}$, the cavity and the qubit density operators remain approximately factorized, such that the cavity behaves as a quasi-ideal source of work inducing a classical drive on the qubit. As a consequence, the qubit follows a unitary evolution ruled by the effective Hamiltonian $\hat{H}^{\text{eff}}_{Q}(t)=\text{Tr}_{C}\{\hat{V}_{\text{QC}}\hat{\rho}_{C}(t)\}=i\hbar g(\left\langle a(t)\right\rangle\sigma_{+}-\langle{a^{\dagger}(t)}\rangle\sigma_{-})$, featuring Rabi oscillations with frequency $\Omega\simeq g|\left\langle a(0)\right\rangle|=g\sqrt{n_{0}}$ (see Fig. 1(c) and [32]). To identify thermodynamic evidence of the work source behavior of the cavity, and deviations from it, we plot in Fig. 2 the work $W_{C}$ and the heat $Q_{C}$ provided by the cavity to the qubit, together with the variation of the cavity’s internal energy $\Delta E_{C}$. We see that for large enough values of $n_{0}$, the cavity mostly provides work and almost no entropy is produced during a single Rabi oscillation. We formalize this result by expanding the exact cavity and qubit state $\hat{\rho}(t)=\hat{U}(t)(\hat{\rho}_{Q}(0)\otimes\hat{\rho}_{C}(0))\hat{U}^{\dagger}(t)$ as power series of $1/n_{0}$. For $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$, the qubit population $P_{e}(t)=\langle e|\hat{\rho}_{Q}(t)|e\rangle$ and coherence $C_{eg}(t)=\langle e|\hat{\rho}_{Q}(t)|g\rangle$ read:
@@
where $\theta=2\Omega t=2gt\sqrt{n_{0}}$ stands for the rotation angle in the Bloch sphere. The complete expression of first order corrections $\delta P_{e}(\theta,\bar{n})/n_{0}$ and $\delta C_{eg}(\theta,\bar{n})/n_{0}$ to an ideal Rabi oscillation are given in the Supp. Mat., and decay as $1/n_{0}$ in the classical limit. These first-order truncations show good agreement with exact numerical results over short time-scales of the order of $\Omega^{-1}$, as shown in Fig. 1(c). In addition, our expansion shows that the amount of exchanged heat $Q_{C}$ and the entropy production scale as $1/n_{0}$ (see inset of Fig. 2), confirming that the cavity can be considered as an ideal work source in this regime.
Figure 2: Heat $Q_{C}$ (red), work $W_{C}$ (green) and internal energy variation of the cavity $\Delta E_{C}$ (blue) as a function of $\theta=2gt\sqrt{n_{0}}$ for $n_{0}=500$ and $\bar{n}=1$, obtained from the second order in $1/n_{0}$. Exact expressions with truncated cavity Hilbert space are displayed in black (not distinguishable). Inset: Heat provided by the cavity during a single Rabi oscillation ($\theta=2\pi$), obtained from the expansion, as a function of $n_{0}$, in semilog scale.
Measurement-induced decoherence – It is known from previous calculations [33], as well as measurements on Rydberg atoms in cavities [34], that the qubit decoheres, and Rabi oscillations collapse, over a time scale $t_{c}\sim g^{-1}$ (see Fig. 1c). This decoherence is due to the evolution of the field towards two orthogonal states, with macroscopically different values of the complex field amplitudes $\left\langle a(t)\right\rangle$ conditioned on the state of the qubit. This evolution can be interpreted as a measurement of the qubit by the cavity  [^1] A more correct terminology here would be a “pre-measurement”. The measurement would be complete only if the measurement outcome is copied in a large number of environmental degrees of freedom [35]. If the cavity and qubit continue to evolve in isolation, revivals of the Rabi oscillations would eventually be observed [33]. However, the measurement analogy remains fully consistent on the time interval considered in this letter.[36].
@@
Autonomous feedback and qubit purification. Remarkably, the behavior of the machine presents a third regime on an even longer timescale $t\sim\sqrt{n_{0}}g^{-1}$, in which the cavity performs an effective drive depending on the result of the measurement in the {|±$\{|\pm_{y}\rangle\}$ basis, thereby performing an autonomous feedback. We show below that this feedback is able to purify any initial qubit state when $n_{0}$ is sufficiently large. We first note that at $t\sim t_{c}$, the cavity field $\text{Tr}\{a\hat{\rho}_{C}^{\nu\nu}\}=\alpha_{\nu}(t)$ depends on the measurement result $\nu$. It is instructive to analyze the joint qubit-cavity dynamics along the two independent branches associated with the qubit found in |±$|\pm_{y}\rangle$. As, inside each branch, the qubit-cavity state is factorized up to ${\cal O}(1/n_{0})$ corrections, one can again consider the cavity as an ideal work source, inducing a classical drive on the qubit of the form $\hat{H}^{\text{eff}(\nu)}_{Q}(t)=i\hbar g(\alpha_{\nu}(t)\sigma_{+}-\alpha_{\nu}^{*}(t)\sigma_{-})$ [32], where, unlike early times ($t<t_{c}$), the effective driving Hamiltonian depends on the measurement result. At $t_{c}$, the two Hamiltonians almost coincide with $\hat{H}_{Q}^{\text{eff}}(0)$, and the qubit measured state is aligned with the field, whatever $\nu$. Then, two different field axes emerge. From Eq. (4), we see that they both lie in the equator and rotate in opposite directions with angular velocities ±$\pm\frac{g}{2\sqrt{n_{0}}}$ much smaller, by a factor of $1/n_{0}$, than the energy splitting of $\hat{H}^{\text{eff}(\nu)}_{Q}(t)$, still given by the Rabi frequency $\Omega$. As a consequence, the qubit follows adiabatically the rotation axis inside each measurement branch (see Fig. 3). The two field axes meet again at $t=t_{\text{min}}=\sqrt{n_{0}}\pi g^{-1}$ where they overlap with the $x$-axis of the Bloch sphere, driving in both branches the qubit to state $|+_{x}\rangle$. As a consequence of this autonomous feedback, an arbitrary initial qubit state $\hat{\rho}_{Q}(0)$ is deterministically purified and brought to $|+_{x}\rangle$ (or generally for a phase $\phi_{0}$ to state $(1/\sqrt{2})\left(|g\rangle+e^{i\phi_{0}}|e\rangle\right)$) at time $t_{\text{min}}$.
@@
It is worth noting that this ideal purification occurs in the limit $n_{0}\to\infty$, such as $t_{\text{min}}$ also diverges. Nevertheless, our expansion at first order in ${\cal O}(1/n_{0})$ allows us to identify corrections for finite $n_{0}$. Mainly, the corrections are due to correlations between the cavity and the qubit inside each branch, which are responsible for a purity loss of the qubit state. As a result, the qubit achieves a non-zero minimum entropy (scaling as ${\cal O}(1/n_{0})$), at a time $t_{\text{min}}$ which slightly differs from $\sqrt{n_{0}}\pi$ (See Fig. 1(a)). In summary, there is a trade-off between the purity of the attained qubit state and the time needed to purify the qubit.
@@
Thermodynamic signature of an autonomous Maxwell demon – We now use our thermodynamic framework to analyze the purification process, and show that it can be interpreted as an autonomous Maxwell demon. We assume that the qubit is initially prepared in the maximally mixed state $\rho_{Q}(0)=\mathds{1}/2$. Fig. 4 shows the evolution of the qubit and cavity entropies, together with their quantum mutual information $I_{CQ}(t)$. We also plotted the work $W_{C}(t)$ and heat $Q_{C}(t)$ provided by the cavity, together with its internal energy variation $\Delta E_{C}$. We first note that the machine is able to decrease the qubit entropy (initially equal to its maximal value $\log 2$) down to a minimum value reached at time $t_{\text{min}}$, at the expanse of work performed by the cavity, akin to a Szilard engine [38, 39].
@@
Supplemental Material: The Jaynes Cummings model as an autonomous Maxwell demon
@@
with $\omega_{0}$ the common transition frequency and $g$ the coupling constant between the qubit and the cavity. The operators $a$ and $a^{\dagger}$ are the usual bosonic annihilation and creation operators and $\sigma_{+}=|e\rangle\langle g|=\sigma_{-}^{\dagger}$ with $|e\rangle$ and $|g\rangle$ the excited and ground states of the two-level system, respectively. The total unitary evolution operator can be written as follows [33]:

### v2

# The Jaynes-Cummings model as an autonomous Maxwell demon

@@ Abstract
We revisit the Jaynes-Cummings model as an autonomous thermodynamic machine, where a qubit is driven by a cavity containing initially a large coherent field. Our analysis reveals a transition between the expected behavior of ideal-work source of the cavity at short times, and a long-time dynamics where the cavity autonomously measures the qubit and exerts a result-dependent drive. This autonomous feedback then purifies the qubit irrespective of its initial state. We show that the cavity functions thermodynamically as an autonomous Maxwell demon, trading mutual information for cooling power.
Autonomous machines, performing tasks without external control, are an important concept in thermodynamics [1, 2]. Their time evolution derive solely from their initial non-equilibrium condition, akin to automaton. This concept is particularly important in the case of quantum machines [3, 4, 5, 6, 7, 8, 9], where the coherence of microscopic constituents is crucial, as avoiding classical controllers empirically mitigates decoherence. This motivates research into autonomous (or coherent) feedback [10, 11, 12, 13, 14, 15], where a quantum system is able to drive a target system conditionally on the values of a few of its observables, e.g. applied to quantum error correction [16, 17, 18]. When these controllers perform thermodynamic tasks, such as work extraction or cooling, they are called autonomous Maxwell demons [19, 20, 21, 22, 23, 24, 25]. A thorough characterization of these machines is of major fundamental interest, as it is linked to the long-standing question of how the dynamics induced by quantum measurements [26, 27, 28] and the laws of thermodynamics [29, 30, 31] emerge in an entirely quantum framework.
@@
Model – We consider a qubit $Q$ (e.g. a two-level atom) interacting with a harmonic oscillator (hereafter dubbed the cavity $C$) according to the Jaynes-Cummings interaction $\hat{V}_{\text{QC}}=i\hbar g(a\sigma_{+}-a^{\dagger}\sigma_{-})$, where $a$ denotes the bosonic annihilation operator of the cavity, and $\sigma_{-}=|g\rangle\langle e|=\sigma_{+}^{\dagger}$ the lowering operator of the qubit. We focus on resonant conditions, where the qubit and cavity Hamiltonians are $\hat{H}_{Q}=\hbar\omega_{0}\sigma_{+}\sigma_{-}$ and $\hat{H}_{C}=\hbar\omega_{0}a^{\dagger}a$, respectively. We assume that the cavity is initially prepared in a coherently-displaced thermal state $\hat{\rho}_{C}(0)=D(\alpha_{0})\hat{w}_{\beta(0)}D^{\dagger}(\alpha_{0})$, where $\hat{w}_{\beta(0)}=\frac{e^{-\beta(0)\hbar\omega_{0}a^{\dagger}a}}{Z_{C}(0)}$ is a thermal equilibrium state at inverse temperature $\beta(0)$, with $Z_{C}(0)=\text{Tr}\{e^{-\beta(0)\hbar\omega_{0}a^{\dagger}a}\}$. $D(\alpha_{0})=e^{\alpha_{0}a^{\dagger}-\alpha_{0}^{*}a}$ denotes the phase-space displacement operator with amplitude $\alpha_{0}=\sqrt{n_{0}}e^{i\phi_{0}}$. We are interested in the limit $n_{0}\gg\bar{n},1$, with $\bar{n}=(e^{\beta(0)\hbar\omega_{0}}-1)^{-1}$ the initial thermal occupation of the cavity, in which the cavity is expected to induce a classical drive for the qubit.
Figure 1: (a) Entropy variation $\Delta S_{Q}$ of the qubit over a long time scale for different initial qubit states $|\psi_{1}\rangle=|e\rangle\langle e|$, $|\psi_{2}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+e^{i\frac{\pi}{4}}|g\rangle)$, and $|\psi_{3}\rangle=|+_{y}\rangle=\frac{1}{\sqrt{2}}(|e\rangle+i|g\rangle)$. (b) zoom in the early times of (a), in semilog scale. (c) Excited state population $P_{e}=\langle e|\hat{\rho}_{Q}(t)|e\rangle$ (in blue) and coherence $C_{eg}=\langle e|\hat{\rho}_{Q}(t)|g\rangle$ (in orange and red) of the qubit when it is initialized in $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$. The dots correspond to expansions up to second order in $1/n_{0}$ and the solid lines are obtained from the numerical evaluation of the exact expressions, truncating the cavity Hilbert space to $N_{\text{ph}}=500$. Parameters: $n_{0}=100$, $\bar{n}=1$.
@@
We find that the dynamics of the autonomous machine features three successive regimes with very different thermodynamic interpretations, that can be identified from the time-evolution of the qubit’s von Neumann entropy as shown in Fig. 1(a) and (b). We analyze them in the following.
Unitary regime – For times $t\ll g^{-1}$, the cavity and the qubit density operators remain approximately factorized, such that the cavity behaves as a quasi-ideal source of work inducing a classical drive on the qubit. As a consequence, the qubit follows a unitary evolution ruled by the effective Hamiltonian $\hat{H}^{\text{eff}}_{Q}(t)=\text{Tr}_{C}\{\hat{V}_{\text{QC}}\hat{\rho}_{C}(t)\}=i\hbar g(\left\langle a(t)\right\rangle\sigma_{+}-\langle{a^{\dagger}(t)}\rangle\sigma_{-})$, featuring Rabi oscillations with frequency $\Omega\simeq g|\left\langle a(0)\right\rangle|=g\sqrt{n_{0}}$ (see Fig. 1(c) and [32]). To identify thermodynamic evidence of the work source behavior of the cavity, and deviations from it, we plot in Fig. 2 the work $W_{C}$ and the heat $Q_{C}$ provided by the cavity to the qubit, together with the variation of the cavity’s internal energy $\Delta E_{C}$. We see that for large enough values of $n_{0}$, the cavity mostly provides work and almost no entropy is produced during a single Rabi oscillation. We formalize this result by expanding as a power series of $1/n_{0}$ the exact cavity and qubit state in the interaction picture $\hat{\rho}(t)=\hat{U}(t)(\hat{\rho}_{Q}(0)\otimes\hat{\rho}_{C}(0))\hat{U}^{\dagger}(t)$, where the analytical expression of $\hat{U}(t)$ is provided in [32]. For $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$, the qubit population $P_{e}(t)=\langle e|\hat{\rho}_{Q}(t)|e\rangle$ and coherence $C_{eg}(t)=\langle e|\hat{\rho}_{Q}(t)|g\rangle$ read:
@@
where $\theta=2\Omega t=2gt\sqrt{n_{0}}$ stands for the rotation angle in the Bloch sphere. The complete expression of first order corrections $\delta P_{e}(\theta,\bar{n})/n_{0}$ and $\delta C_{eg}(\theta,\bar{n})/n_{0}$ to an ideal Rabi oscillation are given in [32], and decay as $1/n_{0}$ in the classical limit. These first-order truncations show good agreement with exact numerical results over short time-scales of the order of $\Omega^{-1}$, as shown in Fig. 1(c). In addition, our expansion shows that the amount of exchanged heat $Q_{C}$ and the entropy production scale as $1/n_{0}$ (see Fig. 2), confirming that the cavity can be considered as an ideal work source in this regime.
Figure 2: Heat $Q_{C}$ (red), work $W_{C}$ (green) and internal energy variation of the cavity $\Delta E_{C}$ (blue) as a function of $\theta=2gt\sqrt{n_{0}}$, for $\hat{\rho}_{Q}(0)=|e\rangle\langle e|$, $n_{0}=500$ and $\bar{n}=1$. Dots are obtained from the second order expansion in $1/n_{0}$, and exact expressions with truncated cavity Hilbert space are displayed in solid black (not distinguishable). For smaller values of $n_{0}$, the heat provided by the cavity is non-negligible, as shown by the dashed black line for $n_{0}=10$.
Measurement-induced decoherence – It is known from previous studies [33, 34, 35], as well as measurements on Rydberg atoms in cavities [36, 37], that the qubit decoheres, and Rabi oscillations collapse, over a time scale $t_{c}\sim g^{-1}$ (see Fig. 1(c)). This decoherence is due to the evolution of the field towards two orthogonal states, with macroscopically different values of the complex field amplitudes $\left\langle a(t)\right\rangle$ conditioned on the state of the qubit. This evolution can be interpreted as a measurement of the qubit by the cavity  [^1] A more correct terminology here would be a “pre-measurement”. The measurement would be complete only if the measurement outcome is copied in a large number of environmental degrees of freedom [38]. If the cavity and qubit continue to evolve in isolation, revivals of the Rabi oscillations would eventually be observed [35]. However, the measurement analogy remains fully consistent on the time interval considered in this letter.[39, 40].
@@
Autonomous feedback and qubit purification. Remarkably, the behavior of the machine presents a third regime on an even longer timescale $t\sim\sqrt{n_{0}}g^{-1}$, in which the cavity performs an effective drive depending on the result of the measurement in the {|±$\{|\pm_{y}\rangle\}$ basis, thereby performing an autonomous feedback. This feedback is able to purify any initial qubit state when $n_{0}$ is sufficiently large [40]. We first note that at $t\sim t_{c}$, the cavity field $\text{Tr}\{a\hat{\rho}_{C}^{\nu\nu}\}=\alpha_{\nu}(t)$ depends on the measurement result $\nu$. It is instructive to analyze the joint qubit-cavity dynamics along the two independent branches associated with the qubit found in |±$|\pm_{y}\rangle$. As, inside each branch, the qubit-cavity state is factorized up to ${\cal O}(1/n_{0})$ corrections, one can again consider the cavity as an ideal work source, inducing a classical drive on the qubit of the form $\hat{H}^{\text{eff}(\nu)}_{Q}(t)=i\hbar g(\alpha_{\nu}(t)\sigma_{+}-\alpha_{\nu}^{*}(t)\sigma_{-})$ [32], where, unlike early times ($t<t_{c}$), the effective driving Hamiltonian depends on the measurement result. At $t_{c}$, the two Hamiltonians almost coincide with $\hat{H}_{Q}^{\text{eff}}(0)$, and the qubit measured state is aligned with the field, whatever $\nu$. Then, two different field axes emerge. From Eq. (4), we see that they both lie in the equator and rotate in opposite directions with angular velocities ±$\pm\frac{g}{2\sqrt{n_{0}}}$ much smaller, by a factor of $1/n_{0}$, than the energy splitting of $\hat{H}^{\text{eff}(\nu)}_{Q}(t)$, still given by the Rabi frequency $\Omega$. As a consequence, the qubit follows adiabatically the rotation axis inside each measurement branch (see Fig. 3). The two field axes meet again at $t=t_{\text{min}}=\sqrt{n_{0}}\pi g^{-1}$ where they overlap with the $x$-axis of the Bloch sphere, driving in both branches the qubit to state $|+_{x}\rangle$. As a consequence of this autonomous feedback, an arbitrary initial qubit state $\hat{\rho}_{Q}(0)$ is deterministically purified and brought to $|+_{x}\rangle$ (or generally for a phase $\phi_{0}$ to state $(1/\sqrt{2})\left(|g\rangle+e^{i\phi_{0}}|e\rangle\right)$) at time $t_{\text{min}}$.
@@
It is worth noting that this ideal purification occurs in the limit $n_{0}\to\infty$, such that $t_{\text{min}}$ also diverges. Nevertheless, our expansion at first order in ${\cal O}(1/n_{0})$ allows us to identify corrections for finite $n_{0}$. Mainly, the corrections are due to correlations between the cavity and the qubit inside each branch, which are responsible for a purity loss of the qubit state. As a result, the qubit achieves a non-zero minimum entropy (scaling as ${\cal O}(1/n_{0})$), at a time $t_{\text{min}}$ which slightly differs from $\sqrt{n_{0}}\pi g^{-1}$ (See Fig. 1(a)). In summary, there is a trade-off between the purity of the attained qubit state and the time needed to purify the qubit.
@@
Thermodynamic signature of an autonomous Maxwell demon – We now use our thermodynamic framework to analyze the purification process, and show that it can be interpreted as an autonomous Maxwell demon. We assume that the qubit is initially prepared in the maximally mixed state $\rho_{Q}(0)=\mathds{1}/2$. Fig. 4 shows the evolution of the qubit and cavity entropies, together with their quantum mutual information $I_{CQ}(t)=S_{C}(t)+S_{Q}(t)-S_{CQ}(t)$. We also plotted the work $W_{C}(t)$ and heat $Q_{C}(t)$ provided by the cavity, together with its internal energy variation $\Delta E_{C}$. We first note that the machine is able to decrease the qubit entropy (initially equal to its maximal value $\log 2$) down to a minimum value reached at time $t_{\text{min}}$, at the expense of work performed by the cavity, akin to a Szilard engine [42, 43].
@@
Supplemental Material: The Jaynes-Cummings model as an autonomous Maxwell demon
@@
with $\omega_{0}$ the common transition frequency and $g$ the coupling constant between the qubit and the cavity. The operators $a$ and $a^{\dagger}$ are the usual bosonic annihilation and creation operators and $\sigma_{+}=|e\rangle\langle g|=\sigma_{-}^{\dagger}$ with $|e\rangle$ and $|g\rangle$ the excited and ground states of the qubit, respectively.
Due to the resonance condition, the evolution of the total system in the interaction picture is governed by Hamiltonian $\hat{H}_{I}(t)=\hat{V}_{\text{QC}}$. We deduce the interaction picture evolution operator $\hat{U}(t)=e^{-\frac{it}{\hbar}\hat{V}_{QC}}$, which can expanded as [35]: